OpenStudy (anonymous):

solve the proportion 1/2=x/19 A) x=1/38 B) x=19 C) x=38 D) x=9 1/2

OpenStudy (anonymous):

\[\frac{1}{2}=\frac{x}{19}\iff \frac{19}{2}=x\]

OpenStudy (skullpatrol):

Did that^ help?

OpenStudy (anonymous):

nope, i still dont understand

OpenStudy (kamille):

I will try to draw you!